OpenStudy (anonymous):

Calculate the standard cell potential at 25 ∘C for the reaction X(s)+2Y+(aq)→X2+(aq)+2Y(s) where ΔH∘ = -703kJ and ΔS∘ = -245J/K .

OpenStudy (aaronq):

Combine: \(\Delta G=\Delta H-T\Delta S\) and \(\Delta G=-n\xi F\) \(\Delta H-T\Delta S=-n\xi^o F\) \(\xi ^o=\dfrac{-(\Delta H-T\Delta S)}{nF}=\dfrac{T\Delta S-\Delta H}{nF}\) \(\xi ^o\) =Standard potential \(F\)=faradays constant n=stoichiometry of electrons transferred (equal to 2 here)